# ChessGPT: Bridging Policy Learning and Language Modeling

### v1

@@ 5.3 Policy evaluation
Checkmate in one We incorporate the checkmate-in-one task from Big-Bench (srivastava2022beyond, ) into our evaluation methods. This task is designed to challenge the model’s ability to identify a move in a given PGN that would result in a checkmate. By doing so, it measures the model’s capacity to comprehend and apply the rules of chess. The model is essentially required to discern a move that not only places the opponent’s king under attack but also ensures that the king cannot evade capture in the next move.
@@
Further analyzing the performance of different models across varying Elo Ratings is crucial for understanding the observed results. The minor variations in move scores for different Elo Rating scenarios in table 8 indicate that ChessGPT-Base may struggle to effectively incorporate Elo Rating information into its decision-making process. This could be due to the model’s limited understanding of the nuanced characteristics associated with distinct Elo Ratings. The complexity of the task and the challenges in accurately accounting for diverse playing styles further contribute to the limited variations in move scores across different Elo Ratings. Consequently, neglecting this information can lead to the model learning an average policy for each Elo Rating, resulting in subpar overall performance. Similar findings were observed in the black chess test, and to further validate this viewpoint, we conducted input attention visualization. Refer to appendix D.1.2 for more details.
@@ 5.4 Qualitative results
We also perform qualitative comparison between our models (ChessGPT-Chat and ChessGPT-Base) and the baselines. We ask the language models a series of questions ranging from factual knowledge of chess as well as requesting the models to perform some operational tasks related to chess. We found that ChessGPT-base performed similarly to RedPajama: both models can sometimes produce factual answers for some of the questions but they failed to generate coherent answers when asked to perform tasks such as providing commentary on chess moves or converting the PGN notation to FEN. ChessGPT-Chat gives more factual answers and demonstrates better performance when prompted to generate analysis and perform other chess-related tasks. Refer to appendix E for qualitative analysis.
@@ 6 Conclusion
In this paper, we introduce a new large-scale dataset and benchmark on chess to encourage of study of the interplay between historical policy data and natural language knowledge. We accompany our dataset with an evaluation framework for assessing language models’ capability in chess. We showcase two models, ChessCLIP and ChessGPT, that demonstrate promising results for learning the interplay between language and action. Nevertheless, our results indicate that we are only beginning to understand how to bridge the gap between policy learning and language modeling and we discuss more on the future directions of our dataset in appendix F. We hope that our dataset and benchmark can make future research on policy and language alignment more accessible.
Checklist
The checklist follows the references. Please read the checklist guidelines carefully for information on how to answer these questions. For each question, change the default [TODO] to [Yes] , [No] , or [N/A] . You are strongly encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description. For example:
Did you include the license to the code and datasets? [Yes] See Section 2.
Did you include the license to the code and datasets? [No] The code and the data are proprietary.
Did you include the license to the code and datasets? [N/A]
Please do not modify the questions and only use the provided macros for your answers. Note that the Checklist section does not count towards the page limit. In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes] We describe it in appendix G
Did you discuss any potential negative societal impacts of your work? [Yes] We describe it in appendix G
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We open source our code, model, and dataset at https://github.com/waterhorse1/ChessGPT.
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Refer to section D.1.1 and section D.1.2
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] Refer to the table and we report the standard deviation.
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Refer to section D.1.3
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes]
Did you mention the license of the assets? [Yes]
Did you include any new assets either in the supplemental material or as a URL? [Yes] We open source our code, model, and dataset at https://github.com/waterhorse1/ChessGPT.
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Refer to appendix C
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] Refer to appendix C
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
[appendix]
@@ B.4 Forsyth–Edwards Notation (FEN)
Appendix C Data Collection and Preprocessing
C.1 Online Chess websites
@@
We utilize Beautifulsoup111https://www.crummy.com/software/BeautifulSoup/and Playwright [1] to parse HTML pages and locate all the texts and PGNs. We further transfer those PGNs into text which helps us build mixed game-language datasets from these sources. We record crucial details such as the URL, author, title, time, and the actual content of the articles.
C.2 Online Chess forums
We choose 5 chess forums and follow basically similar way with appendix C.1 to scrape forum text data.
C.3 Annotated PGN
@@
Lichess studies Lichess Study provides a rich collection of annotated PGNs. The annotations are embedded in PGNs, explaining the insight of the moves. Users can conveniently search for studies based on keywords or specific topics like Sicilian Defense, Puzzles, or d5. To enhance the searching process, we collect a comprehensive set of 54 popular keywords. Our implementation leverages Selenium’s [43] built-in functions to efficiently parse HTML pages and simulate the searching process. Additionally, we use Lichess APIs222https://lichess.org/apito request for PGN games associated with a specific study ID.
Chess publishing This contains commercial annotated PGNs from Chesspublishing.com333https://www.chesspublishing.com/content/, so we do not open source this source of data.
Megabase This contains commercial annotated PGNs from Megabase2023444https://shop.chessbase.com/en/products/mega_database_2023, so we do not open source this source of data.
Pgnlib We collect annotated PGN from Pgnlib555https://www.angelfire.com/games3/smartbridge/.
Pathtochessmastery We collect annotated PGNs from Path to Chess Mastery666https://www.pathtochessmastery.com/.
Gameknot We use Selenium [43] to scrape annotated PGNs from gameknot777https://gameknot.com/list_annotated.pl?u=all.
C.4 Existing Datasets
We mainly extract all chess-related language corpus from existing dataset C4 [39], Pile [18], Oscar [33], Wikipedia [17] and RedPajama [48]. To extract chess-related language corpus, we first filter language corpus that contains the keyword ’chess’. We further utilize the deberta-v3-base-tasksource-nli model888https://huggingface.co/sileod/deberta-v3-base-tasksource-nli, which can conduct zero-shot classification based on classes specified by users. We set two classes: chess-related and non-chess-related for this model. We input the first 2000 characters of each text example to the model and set the threshold as 0.5.
C.5 YouTube transcripts dataset
To collect the youtube transcripts dataset, we first gather 19 keywords and utilize scrapetube999https://github.com/dermasmid/scrapetube to gather all related channels. We extract all videos from these channels and use the same deberta-v3-base-tasksource-nli model mentioned in appendix C.4 to filter all video transcripts and also the corresponding videos that are not relevant to chess. It is fairly easy to extract transcripts from videos and the main difficulty is to extract the FEN chessboard representations from videos. Here we mainly utilize two steps to extract specific FEN from chess videos.
C.5.1 Extract Chess Board from videos
@@
C.5.2 Convert Chess board image to FEN text format
@@
Chess board decomposition The aim of this section is to breakdown a whole chess board into 64 small tiles (8 rows * 8 columns), where each tile contains only one chess piece. To achieve this, we initially convert the RGB image to grayscale, preparing for the line detection process. Subsequently, we make two convolutional kernels to find horizontal and vertical gradients101010https://github.com/Elucidation/tensorflow_chessbot/blob/master/tensorflow_compvision.ipynb. The Hough Transform is then applied to detect vertical and horizontal lines and filter out seven vertical and seven horizontal lines that fit the demand. Finally, we divide the board into 64 tiles by having the position of the 14 lines.
Piece classification To facilitate model training and evaluation, we employ an open source chess image dataset on Kaggle111111https://www.kaggle.com/datasets/koryakinp/chess-positions which contains 80k training images and 20k testing images. Each tile can be classified into one of the 13 categories (p, r, b, n, k, q, P, R, B, N, K, Q, and space) which is detailed in Appendix B.3. We implement a model in PyTorch which uses a pre-trained ResNet18 [19] due to its well-established performance in image classification tasks. To adapt the model to our specific problem, we replaced the original fully connected layer with a new layer consisting of 13 output neurons, corresponding to the 13 pieces categories. We train the model on 40000 images with hyperparameters shown in Table 9.
@@ C.6 Lichess dataset
C.7 Pro-player dataset
We collect our pro-player dataset from PGN Mentor121212https://www.pgnmentor.com/.
C.8 Chess books
@@
C.9 CCRL
We collect our CCRL dataset without comments from the official website131313https://ccrl.chessdom.com/ccrl/4040/ for three settings of time control: CCRL BLITZ, CCRL 40/2 FRC and CCRL 40/15.
C.10 Chess puzzles
We collect our chess puzzles from the Lichess puzzle dataset141414https://database.lichess.org/#puzzles.
C.11 Chess modeling data
@@ C.12 Preprocessing
StackExchange We use the forums Chess StackExchange151515https://chess.stackexchange.com/. We preprocess the StackExchange data in the same way as done in RedPajama. Concretely, for each question, we first clean up the text via simple regular expressions and remove some HTML tags in the text. Then we prefix the question with Q: and the answers with A: and then concatenate the questions with all answers.
@@
Annotated PGNs We conduct language filtering using Fasttext161616https://fasttext.cc/docs/en/language-identification.html in ChessCLIP preprocessing. And we conduct the same preprocessing as we do in Lichess database in ChessGPT training preprocessing.
@@
C.13 Licenses and dataset cards
@@
Appendix D Implementation and Evaluation Details
@@
D.1 Implmenetation details
D.1.1 ChessCLIP
@@
For the ChessCLIP model, we instantiate a ChessCLIP model with a pair of text encoder and a board/action encoder. For board/action encoder, we use a ResNet [19] architecture that conditions the action encoding via a modified FiLM layer [36]. We encode the board positions and moves using the same scheme as those used by Leela Chess Zero (lc0) [24], which is similar to the encoding used by AlphaZero [42] for encoding positions and moves in chess. Concretely, the board positions are encoded as a feature map and the actions are encoded as a vector. For the text encoder, we follow the same architecture as with the original OpenAI CLIP model and we only fine-tune the last two layers of pretrained OpenAI text encoder. Our implementation is based on the open-source implementation 171717https://github.com/mlfoundations/open_clip of CLIP. We show our training hyper-parameters in table 11.
@@ D.1.2 ChessGPT
Our base-training code refers to llama-finetune181818https://github.com/chaoyi-wu/Finetune_LLAMA and our sft-training code refers to the Alpaca [47] and Fastchat [12]. The training hyperparameters for ChessGPT-Base and ChessGPT-Chat are shown in table 12 and table 13.

### v2

@@ 5.3 Policy evaluation
Checkmate in one We incorporate the checkmate-in-one task from Big-Bench srivastava2022beyond into our evaluation methods. This task is designed to challenge the model’s ability to identify a move in a given PGN that would result in a checkmate. By doing so, it measures the model’s capacity to comprehend and apply the rules of chess. The model is essentially required to discern a move that not only places the opponent’s king under attack but also ensures that the king cannot evade capture in the next move.
@@
Further analyzing the performance of different models across varying Elo Ratings is crucial for understanding the observed results. The minor variations in move scores for different Elo Rating scenarios in Table 8 indicate that ChessGPT-Base may struggle to effectively incorporate Elo Rating information into its decision-making process. This could be due to the model’s limited understanding of the nuanced characteristics associated with distinct Elo Ratings. The complexity of the task and the challenges in accurately accounting for diverse playing styles further contribute to the limited variations in move scores across different Elo Ratings. Consequently, neglecting this information can lead to the model learning an average policy for each Elo Rating, resulting in subpar overall performance. Similar findings were observed in the black chess test, and to further validate this viewpoint, we conducted input attention visualization. Refer to Section F.1.2 for more details.
@@ 5.4 Qualitative results
We also perform qualitative comparison between our models (ChessGPT-Chat and ChessGPT-Base) and the baselines. We ask the language models a series of questions ranging from factual knowledge of chess as well as requesting the models to perform some operational tasks related to chess. We found that ChessGPT-base performed similarly to RedPajama: both models can sometimes produce factual answers for some of the questions but they failed to generate coherent answers when asked to perform tasks such as providing commentary on chess moves or converting the PGN notation to FEN. ChessGPT-Chat gives more factual answers and demonstrates better performance when prompted to generate analysis and perform other chess-related tasks. Refer to Appendix H for qualitative analysis.
@@ 6 Conclusion
In this paper, we introduce a new large-scale dataset and benchmark on chess to encourage of study of the interplay between historical policy data and natural language knowledge. We accompany our dataset with an evaluation framework for assessing language models’ capability in chess. We showcase two models, ChessCLIP and ChessGPT, that demonstrate promising results for learning the interplay between language and action. Nevertheless, our results indicate that we are only beginning to understand how to bridge the gap between policy learning and language modeling and we discuss more on the future directions of our dataset in Appendix J. We hope that our dataset and benchmark can make future research on policy and language alignment more accessible.
@@ B.4 Forsyth–Edwards Notation (FEN)
Appendix C Image sources
Here we provide the sources for images used in our Figure 1.
Online game: https://lichess.org/tv
Pro game: https://www.insidethegames.biz/articles/1110959/magnus-carlsen-chess-world-cup-duda
Computer game: https://www.chess.com/news/view/updated-alphazero-crushes-stockfish-in-new-1-000-game-match
Modeling: https://www.chess.com/terms/fen-chess
Puzzle: https://chesspuzzle.net/
General text: https://www.bbc.co.uk/newsround/66233770
Wiki: https://en.wikipedia.org/wiki/Chess
Forum: https://www.chess.com/forum
Blog: https://www.chess.com/blogs
Book: https://www.amazon.co.uk/Chess-Strategy-Action-John-Watson/dp/1901983692
Annotated game: https://www.chess.com/forum/view/game-analysis/kingside-pawn-rush
Youtube video: https://www.youtube.com/watch?v=NAIQyoPcjNM
Reddit: https://www.reddit.com/r/chess/comments/15s65hb/white_to_play_and_reach_2000_chesscom/jwdqpze/?context=3
Value judgement: https://www.thenewatlantis.com/wp-content/uploads/legacy/20190820_TNA58Wilkenfeldbanner.jpg
Appendix D Dataset analysis
| Component | Raw size | Document count | Subset type |
| Lichess | 19.9 GB | 17.5 M | Game |
| Pro-player | 0.37 GB | 0.44 M | Game |
| CCRL | 3.60 GB | 3.00 M | Game |
| Chess puzzles | 1.53 GB | 3.19 M | Game |
| Chess modeling | 0.94 GB | 1.90 M | Game |
| C4-Chess | 0.59 GB | 0.18 M | Language |
| Pile-Chess | 1.10 GB | 0.10 M | Language |
| RedPajama-Chess | 5.65 GB | 0.52 M | Language |
| Oscar-Chess | 3.13 GB | 0.33 M | Language |
| WikiPedia-Chess | 40.3 MB | 11.4 K | Language |
| Chess Blogs | 0.59 GB | 73.2 K | Language |
| Chess Books | 1.86 GB | 8.36 K | Language |
| Chess Forums | 1.05 GB | 0.14 M | Language |
| Annotated Chess Games | 0.66 GB | 245 K | Mixed |
| Youtube transcripts | 0.98 GB | 83.0K | Mixed |
| GPT-4-Chess | 0.95 MB | 3.91 K | Conversation |
| 0.86 GB | 0.41 M | Conversation | |
| Overall | 42.8 GB | 28.1 M | N/A |
D.1 Dataset statistics and metrics
| dataset | num. words | char. rep. | word rep. | special char | stopwords | perplexity |
| ratio | ratio | ratio | ratio | |||
| Chess puzzles | 49.6618 | 0.0104 | 0.0000 | 0.3246 | 0.4771 | 988.3734 |
| Oscar-Chess | 1441.4341 | 0.0596 | 0.0499 | 0.2242 | 0.4119 | 665.6499 |
| Pile-Chess | 2105.2454 | 0.0626 | 0.0205 | 0.2409 | 0.4227 | 497.3883 |
| RedPajama-Chess | 1581.5825 | 0.0532 | 0.0163 | 0.2273 | 0.4218 | 410.1236 |
| StackExchange-Chess | 578.3733 | 0.0591 | 0.0816 | 0.2617 | 0.4835 | 520.1257 |
| Wikipedia-Chess | 463.4980 | 0.0876 | 0.0052 | 0.2604 | 0.3079 | 236.9671 |
| C4-Chess | 510.6041 | 0.0479 | 0.0082 | 0.2131 | 0.4418 | 548.9471 |
InTable 9, we present the dataset statistics breakdown for each data subset, including its raw size, document count, and subset type.
Table 10 shows the properties of the chess-specific language dataset that we use for training ChessGPT. For these datasets, we computed the average number of words (num. words) per example, character repetition ratio (char. rep. ratio), word repetition ratio (word. rep. ratio.), special character ratio (special char. ratio), stopwords ratio and perplexity of the first shard for each subset in the language dataset. These metrics are based on some of the criteria employed by Bloom [25] in their pre-processing pipelines.
D.2 Dataset examples
In this subsection, we offer dataset examples for each subset to help the readers to understand its contents more intuitively and clearly.
Appendix E Data Collection and Preprocessing
E.1 Online Chess websites
@@
We utilize Beautifulsoup333https://www.crummy.com/software/BeautifulSoup/and Playwright [1] to parse HTML pages and locate all the texts and PGNs. We further transfer those PGNs into text which helps us build mixed game-language datasets from these sources. We record crucial details such as the URL, author, title, time, and the actual content of the articles.
E.2 Online Chess forums
We choose 5 chess forums and follow basically similar way with Section E.1 to scrape forum text data.
E.3 Annotated PGN
@@
Lichess studies Lichess Study provides a rich collection of annotated PGNs. The annotations are embedded in PGNs, explaining the insight of the moves. Users can conveniently search for studies based on keywords or specific topics like Sicilian Defense, Puzzles, or d5. To enhance the searching process, we collect a comprehensive set of 54 popular keywords. Our implementation leverages Selenium’s [48] built-in functions to efficiently parse HTML pages and simulate the searching process. Additionally, we use Lichess APIs444https://lichess.org/apito request for PGN games associated with a specific study ID.
Chess publishing This contains commercial annotated PGNs from Chesspublishing.com555https://www.chesspublishing.com/content/, so we do not open source this source of data.
Megabase This contains commercial annotated PGNs from Megabase2023666https://shop.chessbase.com/en/products/mega_database_2023, so we do not open source this source of data.
Pgnlib We collect annotated PGN from Pgnlib777https://www.angelfire.com/games3/smartbridge/.
Pathtochessmastery We collect annotated PGNs from Path to Chess Mastery888https://www.pathtochessmastery.com/.
Gameknot We use Selenium [48] to scrape annotated PGNs from gameknot999https://gameknot.com/list_annotated.pl?u=all.
E.4 Existing Datasets
We mainly extract all chess-related language corpus from existing dataset C4 [43], Pile [19], Oscar [37], Wikipedia [18] and RedPajama [53]. To extract chess-related language corpus, we first filter language corpus that contains the keyword ’chess’. We further utilize the deberta-v3-base-tasksource-nli model101010https://huggingface.co/sileod/deberta-v3-base-tasksource-nli, which can conduct zero-shot classification based on classes specified by users. We set two classes: chess-related and non-chess-related for this model. We input the first 2000 characters of each text example to the model and set the threshold as 0.5.
E.5 YouTube transcripts dataset
To collect the youtube transcripts dataset, we first gather 19 keywords and utilize scrapetube111111https://github.com/dermasmid/scrapetube to gather all related channels. We extract all videos from these channels and use the same deberta-v3-base-tasksource-nli model mentioned in Section E.4 to filter all video transcripts and also the corresponding videos that are not relevant to chess. It is fairly easy to extract transcripts from videos and the main difficulty is to extract the FEN chessboard representations from videos. Here we mainly utilize two steps to extract specific FEN from chess videos.
E.5.1 Extract Chess Board from videos
@@
E.5.2 Convert Chess board image to FEN text format
@@
Chess board decomposition The aim of this section is to breakdown a whole chess board into 64 small tiles (8 rows * 8 columns), where each tile contains only one chess piece. To achieve this, we initially convert the RGB image to grayscale, preparing for the line detection process. Subsequently, we make two convolutional kernels to find horizontal and vertical gradients121212https://github.com/Elucidation/tensorflow_chessbot/blob/master/tensorflow_compvision.ipynb. The Hough Transform is then applied to detect vertical and horizontal lines and filter out seven vertical and seven horizontal lines that fit the demand. Finally, we divide the board into 64 tiles by having the position of the 14 lines.
Piece classification To facilitate model training and evaluation, we employ an open source chess image dataset on Kaggle131313https://www.kaggle.com/datasets/koryakinp/chess-positions which contains 80k training images and 20k testing images. Each tile can be classified into one of the 13 categories (p, r, b, n, k, q, P, R, B, N, K, Q, and space) which is detailed in Appendix B.3. We implement a model in PyTorch which uses a pre-trained ResNet18 [20] due to its well-established performance in image classification tasks. To adapt the model to our specific problem, we replaced the original fully connected layer with a new layer consisting of 13 output neurons, corresponding to the 13 pieces categories. We train the model on 40000 images with hyperparameters shown in Table 11.
@@ E.6 Lichess dataset
E.7 Pro-player dataset
We collect our pro-player dataset from PGN Mentor141414https://www.pgnmentor.com/.
E.8 Chess books
@@
E.9 CCRL
We collect our CCRL dataset without comments from the official website151515https://ccrl.chessdom.com/ccrl/4040/ for three settings of time control: CCRL BLITZ, CCRL 40/2 FRC and CCRL 40/15.
E.10 Chess puzzles
We collect our chess puzzles from the Lichess puzzle dataset161616https://database.lichess.org/#puzzles.
E.11 Chess modeling data
@@ E.12 Preprocessing
StackExchange We use the forums Chess StackExchange171717https://chess.stackexchange.com/. We preprocess the StackExchange data in the same way as done in RedPajama. Concretely, for each question, we first clean up the text via simple regular expressions and remove some HTML tags in the text. Then we prefix the question with Q: and the answers with A: and then concatenate the questions with all answers.
@@
Annotated PGNs We conduct language filtering using Fasttext181818https://fasttext.cc/docs/en/language-identification.html in ChessCLIP preprocessing. And we conduct the same preprocessing as we do in Lichess database in ChessGPT training preprocessing.
@@
E.13 Licenses and dataset cards
@@
Appendix F Implementation and Evaluation Details
@@
F.1 Implmenetation details
F.1.1 ChessCLIP
@@
For the ChessCLIP model, we instantiate a ChessCLIP model with a pair of text encoder and a board/action encoder. For board/action encoder, we use a ResNet [20] architecture that conditions the action encoding via a modified FiLM layer [40]. We encode the board positions and moves using the same scheme as those used by Leela Chess Zero (lc0) [26], which is similar to the encoding used by AlphaZero [47] for encoding positions and moves in chess. Concretely, the board positions are encoded as a feature map and the actions are encoded as a vector. For the text encoder, we follow the same architecture as with the original OpenAI CLIP model and we only fine-tune the last two layers of pretrained OpenAI text encoder. Our implementation is based on the open-source implementation 191919https://github.com/mlfoundations/open_clip of CLIP. We show our training hyper-parameters in Table 13.
@@ F.1.2 ChessGPT
Our base-training code refers to llama-finetune202020https://github.com/chaoyi-wu/Finetune_LLAMA and our sft-training code refers to the Alpaca [52] and Fastchat [12]. The training hyperparameters for ChessGPT-Base and ChessGPT-Chat are shown in Table 14 and Table 15.